OpenStudy (anonymous):

Let x be the width and y be the height of a norman window. Thus, the semi-circle has radius x/2. We must maximize the area of the window, A = xy + pi/2(x/2)^2. The perimeter of the window is 20 = 2y + x + pi(x/2), and so y = __________________ - x/2 - (pi x)/__________________________. Help filling the blanks? I assume I have to solve it a certain way, but I have no idea. Thanks.

OpenStudy (anonymous):

Find the first and second derivative to find the maximum area. \[\huge A = xy + \frac{\pi}{2}(\frac{x^2}{4})\] \[\huge A=xy +\frac{\pi}{8}x^2\] \[\large y=\frac{20-x-\frac{\pi}{2}x}{2}\] Therefore: \[\huge A=\frac{x[20-x-\frac{\pi}{2}x]}{2}+\frac{\pi}{8}x^2\] \[\huge =\frac{20x-x^2-\frac{\pi}{2}x^2}{2}+\frac{\pi}{8}x^2\] \[\huge =\frac{80x-4x^2-2\pi x^2+\pi x^2}{8}\] \[\huge =\frac{80x-4x^2-\pi x^2}{8}\] \[\large \frac{dA}{dx}=\frac{40-4x-\pi x}{4}\] \[\large \frac{d^2A}{dx^2}=\frac{-4-\pi}{4}\] \[\large =-\frac{4+\pi}{4}\] At S.P's dA/dx=0 \[40-4x-\pi x=0\] \[40=4x+\pi x\] \[x(4+\pi)=40\] \[x=\frac{40}{4+\pi}\]

OpenStudy (anonymous):

Sub in x into the equation \[A=\frac{80x-4x^2-\pi x^2}{8}\] To find the maximum Area.

OpenStudy (anonymous):

gotcha thanks! But I don't know what they want me to put in the blanks!

OpenStudy (anonymous):

That's the most simple thing I've ever seen. Never seen someone with this calibre of maths learning this method. Since you have the given equation of the perimeter \[\huge 20 = 2y + x + \pi(\frac{x}{2})\] Rearrange the equation to make y the subject. \[\huge 2y=20-x-\pi(\frac{x}{2})\] \[\huge y=\frac{20-x-\frac{\pi x}{2}}{2}\] Separate the numerator: \[\huge y=\frac{20}{2}-\frac{x}{2}-\frac{\frac{\pi x}{2}}{2}\] SImplify: \[\huge y=10-\frac{x}{2}-[\frac{\pi x}{2}\times \frac{1}{2}]\] \[\huge =10-\frac{x}{2}-\frac{\pi x}{4}\]

OpenStudy (anonymous):

Thank you!